How to calculate the Multiplicative inverse modulo.
like how to calculate 7inverse mod (20). It will be helpful, if the solution or the method is calculative

Answer by stanbon(75887)

How to calculate the Multiplicative inverse modulo.
like how to calculate 7 inverse mod (20).
---
1/7 = (1+20x)/7 = 3 when x = 1
----
1/7 = 3 (mod 20)
====================
Cheers,
Stan H.
============